Integration by parts

1. Integral of 2x tan^-1 (x) dx

2. Integral of x^2 (x+4)^5 dx
1
Expert's answer
2019-12-11T11:05:14-0500

"1. \int2x arctan(x)dx= 2 \int x arctan(x) dx = \begin{vmatrix} u = arctan(x) \\
 dv = xdx\\
du = \frac 1{x^2+1}dx\\
v = \frac {x^2}2 
\end{vmatrix} ="

"=x^2arctan(x) - \int \frac {x^2}{x^2+1}dx =x^2arctan(x) - \int (1 - \frac {1}{x^2+1})dx =\\=x^2arctan(x) - \int1dx+\int\frac{1}{x^2+1} dx =\\"

"= x^2arctan(x)+arctan(x)-x +C"

"2. \int x^2(x+4)^5dx = \int (x^7+20x^6+160x^5+640x^4+1280x^3+1024x^2)dx="

"=\int x^7dx 20\int x^6 dx+160\int x^5dx +640 \int x^4 dx+1280 \int x^3dx+1024\int x^2dx=\\""\\= \frac {x^8}8 +\frac {20x^7}7 + \frac {80x^6}3 + 128x^5 + 320x^4 + \frac{1024x^3}3 +C"